a car is traveling on a straight road at 19 m/s. the driver sees a bird standing in the road 38 meters ahead and applies the brakes, giving constant deceleration of 5 m/s2.

(a) how far away is the driver from the bird when he stops? ( assuming the bird doesn't move)

(b) how long does it take the driver to stop?

(c) the driver was in a 40 miles per hour zone. given that there are approximately 1.6 km in 1 mile, was the driver speeding?




1
Expert's answer
2021-11-24T19:11:03-0500

Solution.

"v_0=19 m\/s;"

"S_1=38m;"

"a=-5 m\/s^2;"

"a) S_2=\dfrac{v^2-v_0^2}{2a};"

"S_2=\dfrac{-19^2}{-10}=36.1 m;"

"S=S_1-S_2;\nS=38-36.1=1.9m;"

"b) a=\dfrac{v-v_0}{t}\implies t=\dfrac{v-v_0}{a};"

"t=\dfrac{-19}{-5}=3.8s;"

"c)" "v=40 mile \/h;"

"1 mile=1.6 km;"

"19 m\/s=68.4 km\/h=42.75 mile\/h;"

Answer: "a)S=1.9 m;"

"b) t=3.8s;"

"c)" the driver exceeded the speed limit.